# Flow Matching for Tabular Data Synthesis

### v1

@@ Appendix A Related Work
Diffusion Models for Tabular Data. Denoising diffusion probabilistic models (DDPMs) have shown great success across domains like vision and language, but their application to tabular data poses unique challenges due to feature heterogeneity and small dataset sizes. TabDDPM (kotelnikov2023tabddpm) addresses this through a dual process: Gaussian diffusion for continuous variables and multinomial diffusion for categorical features. It achieves state-of-the-art (SOTA) utility and privacy performance, outperforming the GAN and VAE baselines. However, its computational cost remains high because of iterative sampling and separate handling of feature types, which may weaken correlation modelling. TabSyn (zhang2024mixedtype) employs a transformer-VAE architecture followed by a diffusion model, while the continuous diffusion model for mixed type tabular data (CDTD) (mueller2025continuous) is a score-matching diffusion model that noises continuous values and categorical embeddings and samples via ODE integration. These two methods improve the utility of tabular data, particularly compared to TabDDPM and other recent SOTAs.
@@
FM and its Variational Extensions. Flow matching (FM) (lipman2023flowmatchinggenerativemodeling; liu2022flowstraightfastlearning; albergo2023stochasticinterpolantsunifyingframework) offers an alternative to diffusion by learning the deterministic dynamics of probability paths. FM has also been developed in terms of latent models for image generation (dao2023flowmatchinglatentspace; schusterbauer2024boosting). In addition, VFM (eijkelboom2024variationalflowmatchinggraph) reformulates FM as a variational inference problem, allowing flow matching through a variational distribution. Consequently, VFM can be implemented to mixed-type data, such as tabular data, through mean-field variational inference. On the other hand, (guzman-cordero2025exponential) extended the VFM specific for the exponential family that also supports mixed-type variables through moment matching and implemented it in tabular data, called TabbyFlow. In addition to tabular, VFM has also been implemented for image generation maticsan2025purrception, controlled molecular generation (eijkelboom2025controlled), non-Euclidean geometry  (zaghen2025variationalflowmatchinggeneral; zaghen2025riemannianvariationalflowmatching), and climate modeling (finn2025generative).
@@
FM Trajectory. Trajectory design is of paramount importance in the field of FM. The paths of optimal transport (OT) are widely preferred because of their straightforward linear interpolation. In contrast, the variance-preserving (VP) path, which takes inspiration from diffusion models, is somewhat underrated (lipman2023flowmatchinggenerativemodeling; liu2022flowstraightfastlearning). Recent studies also introduced different trajectories such as cosine albergo2023buildingnormalizingflowsstochastic, logit-normal labs2025flux1kontextflowmatching. Although the theoretical framework of FM accommodates both types of path, there is a notable gap in empirical evaluation concerning the practical application of different paths, particularly in the realm of tabular data. To address this gap, we present the potential of different paths for tabular data synthesis, despite the fact that OT is the primary implementation.
Appendix B Algorithms
B.1 TabSynFlow
@@ B.2 TabbyFlow
Appendix C Evaluation Details
@@ C.1.1 Ratio of Counts (ROC)
$\mathrm{ROC}_{\text{cell}}=\frac{\min(\mathcal{Y}_{count},\mathcal{Y}^{\prime}_{count})}{\max(\mathcal{Y}_{count},\mathcal{Y}^{\prime}_{count})},$ (C.1)
@@
C.1.2 Confidence Interval Overlap (CIO)
@@
$\mathrm{CIO}_{k}=0.5\left[\frac{\mathrm{overlap}}{u-l}+\frac{\mathrm{overlap}}{u^{\prime}-l^{\prime}}\right],$ (C.2)
where $[l,u]$ and $l^{\prime},u^{\prime}]$ are the confidence intervals from original and synthetic models respectively, and overlap = $\min(u,u^{\prime})-\max(l,l^{\prime})$. A CIO of 1 indicates identical intervals, whereas CIO < 0 indicates no overlap, therefore it is truncated to zero (elliot2023samples). The final CIO score is the mean overlap across all regression coefficients and all fitted models.
C.1.3 Aggregate Utility Score
@@
$\text{Utility}=\frac{\mathrm{ROC_{uni}}+\mathrm{ROC_{biv}}+\mathrm{CIO}}{3}$ (C.3)
@@
C.2 Disclosure Risk Metric
We measure disclosure risk as the probability that an adversary, possessing the quasi-identifier (key variables, $K$), can correctly guess the sensitive value (target $T$) for a record by matching between the synthetic and original data. This approach, adapted from taub2019the, quantifies the additional privacy risk introduced by synthetic data at a per-record granularity, adjusting for the baseline (based on a random draw from $T$), defined by the distribution of $T$ within each equivalence class (combination of $K$) in the original data.
@@ C.2 Disclosure Risk Metric
$TCAP^{\prime}_{j}=P_{\mathcal{D}}(T^{\prime}_{j}|K^{\prime}_{j})=\frac{\sum_{i=1}^{n}\mathbb{I}[T_{i}=T^{\prime}_{j},\,K_{i}=K^{\prime}_{j}]}{\sum_{i=1}^{n}\mathbb{I}[K_{i}=K^{\prime}_{j}]}$ (C.4)
@@
$WEAP_{j}=P_{\mathcal{D}}(T_{j}|K_{j})=\frac{\sum_{i=1}^{n}\mathbb{I}[T_{i}=T_{j},K_{i}=K_{j}]}{\sum_{i=1}^{n}\mathbb{I}[K_{i}=K_{j}]}$ (C.5)
@@
$\text{Risk}_{j}=\max\left\{0,\frac{TCAP^{\prime}_{j}-WEAP_{j}}{1-WEAP_{j}}\right\}$ (C.6)
A value near 0 indicates synthetic data provides no additional risk beyond baseline, while a value near $1$ signals a maximum increase in disclosure risk. Negative values (adversary performs worse than random) are truncated to zero by convention (ran2024multiobjective; elliot2023samples).
Intuition: This metric isolates the extra “leakage” in matching risk—beyond what already exists in the univariate distribution of the target — arising from access to the synthetic data. It thus accounts for singling out, linkage and inference risks (due to the empirical distribution of sensitive values per equivalence class) identified by  A29WP as being critical determinants of risks to anonymity. The disclosure risk for the dataset is the mean of $R_{j}$ over all synthetic records meeting the $\tau$ threshold. The pseudocode to calculate TCAP can be seen in Algorithm 7
@@
11:  Calculate $\mathrm{TCAP}^{\prime}_{j}$ using equation C.4
12:  Calculate $\mathrm{WEAP}_{j}$ using equation C.5
13:  Calculate $\text{Risk}_{j}$ using equation C.6
@@
C.3 Reproducibility
@@
This appendix provides the procedural summary sufficient for manual replication or verification by independent researchers. Table 57 showed detailed variables implemented in the utility and risk evaluation.
@@
Appendix D Additional and Detailed Results
D.1 Variance Comparison Results

### v2

@@ Appendix A Related Work
Diffusion Models for Tabular Data. Denoising diffusion probabilistic models (DDPMs) have shown great success across domains like vision and language, but their application to tabular data poses unique challenges due to feature heterogeneity and small dataset sizes. TabDDPM (Kotelnikov et al. 2023) addresses this through a dual process: Gaussian diffusion for continuous variables and multinomial diffusion for categorical features. It achieves state-of-the-art (SOTA) utility and privacy performance, outperforming the GAN and VAE baselines. However, its computational cost remains high because of iterative sampling and separate handling of feature types, which may weaken correlation modelling. TabSyn (Zhang et al. 2024) employs a transformer-VAE architecture followed by a diffusion model, while the continuous diffusion model for mixed type tabular data (CDTD) (Mueller et al. 2025) is a score-matching diffusion model that noises continuous values and categorical embeddings and samples via ODE integration. These two methods improve the utility of tabular data, particularly compared to TabDDPM and other recent SOTAs.
@@
FM and its Variational Extensions. Flow matching (FM) (Lipman et al. 2023; Liu et al. 2022; Albergo et al. 2023) offers an alternative to diffusion by learning the deterministic dynamics of probability paths. FM has also been developed in terms of latent models for image generation (Dao et al. 2023; Schusterbauer et al. 2024). In addition, VFM (Eijkelboom et al. 2024) reformulates FM as a variational inference problem, allowing flow matching through a variational distribution. Consequently, VFM can be implemented to mixed-type data, such as tabular data, through mean-field variational inference. On the other hand, (Guzmán-Cordero et al. 2025) extended the VFM specific for the exponential family that also supports mixed-type variables through moment matching and implemented it in tabular data, called TabbyFlow. In addition to tabular, VFM has also been implemented for image generation (Matişan et al. 2025), controlled molecular generation (Eijkelboom et al. 2025), non-Euclidean geometry  (Zaghen et al. 2025a; b), and climate modeling (Finn et al. 2025).
@@
FM Trajectory. Trajectory design is of paramount importance in the field of FM. The paths of optimal transport (OT) are widely preferred because of their straightforward linear interpolation. In contrast, the variance-preserving (VP) path, which takes inspiration from diffusion models, is somewhat underrated (Lipman et al. 2023; Liu et al. 2022). Recent studies also introduced different trajectories such as cosine (Albergo and Vanden-Eijnden 2023), logit-normal (Black Forest Labs 2025). Although the theoretical framework of FM accommodates both types of path, there is a notable gap in empirical evaluation concerning the practical application of different paths, particularly in the realm of tabular data. To address this gap, we present the potential of different paths for tabular data synthesis, despite the fact that OT is the primary implementation.
Appendix B TabSyn: Latent Diffusion for Tabular Data
TabSyn (Zhang et al. 2024) represents a state-of-the-art generative approach for mixed-type tabular data. It integrates a transformer-based Variational Autoencoder (VAE) with a score-based diffusion model operating in latent space. The design of TabSyn addresses several key challenges in tabular data synthesis including heterogeneous feature types, complex inter-feature dependencies, and the inefficiency of diffusion processes when employed on high-dimensional structured data.
Within the latent space, noise is introduced through a forward process defined as: $z_{t}=z_{0}+\sigma(t)\varepsilon$. Given linear noise schedule $t$, the corresponding reverse process is an SDE:
$\mathrm{d}z_{t}=-2t\nabla_{z}\log p_{t}(z_{t})\,\mathrm{d}t+\sqrt{2t}\,\mathrm{d}\omega_{t}.$ (B.1)
where $\omega_{t}$ denotes a Wiener process. The model is trained using denoising score matching, with a neural network $\varepsilon_{\theta}(z_{t},t)$ that estimates the scoring function $\nabla_{z_{t}}\log p_{t}(z_{t})=-\varepsilon_{\theta}(z_{t},t)/t$ under linear noise schedule.
TabSyn surpasses traditional GAN and VAE based models as well as the latest diffusion models like TabDDPM and STaSy, across a variety of datasets and evaluation metrics - including unvariate distribution and pairwise correlation preservation, and downstream ML performance. Furthermore, its latent-space formulation allows for missing value imputation via latent inpainting, achieving competitive results in terms of data quality, diversity, and privacy preservation.
Appendix C Algorithms
C.1 TabSynFlow
@@ C.2 TabbyFlow
Appendix D Evaluation Details
@@ D.1.1 Ratio of Counts (ROC)
$\mathrm{ROC}_{\text{cell}}=\frac{\min(\mathcal{Y}_{count},\mathcal{Y}^{\prime}_{count})}{\max(\mathcal{Y}_{count},\mathcal{Y}^{\prime}_{count})},$ (D.1)
@@
D.1.2 Confidence Interval Overlap (CIO)
@@
$\mathrm{CIO}_{k}=0.5\left[\frac{\mathrm{overlap}}{u-l}+\frac{\mathrm{overlap}}{u^{\prime}-l^{\prime}}\right],$ (D.2)
where $[l,u]$ and $l^{\prime},u^{\prime}]$ are the confidence intervals from original and synthetic models respectively, and overlap = $\min(u,u^{\prime})-\max(l,l^{\prime})$. A CIO of 1 indicates identical intervals, whereas CIO < 0 indicates no overlap, therefore it is truncated to zero (Elliot et al. 2023). The final CIO score is the mean overlap across all regression coefficients and all fitted models.
D.1.3 Aggregate Utility Score
@@
$\text{Utility}=\frac{\mathrm{ROC_{uni}}+\mathrm{ROC_{biv}}+\mathrm{CIO}}{3}$ (D.3)
@@
D.2 Disclosure Risk Metric
We measure disclosure risk as the probability that an adversary, possessing the quasi-identifier (key variables, $K$), can correctly guess the sensitive value (target $T$) for a record by matching between the synthetic and original data. This approach, adapted from Taub et al. (2019), quantifies the additional privacy risk introduced by synthetic data at a per-record granularity, adjusting for the baseline (based on a random draw from $T$), defined by the distribution of $T$ within each equivalence class (combination of $K$) in the original data.
@@ D.2 Disclosure Risk Metric
$TCAP^{\prime}_{j}=P_{\mathcal{D}}(T^{\prime}_{j}|K^{\prime}_{j})=\frac{\sum_{i=1}^{n}\mathbb{I}[T_{i}=T^{\prime}_{j},\,K_{i}=K^{\prime}_{j}]}{\sum_{i=1}^{n}\mathbb{I}[K_{i}=K^{\prime}_{j}]}$ (D.4)
@@
$WEAP_{j}=P_{\mathcal{D}}(T_{j}|K_{j})=\frac{\sum_{i=1}^{n}\mathbb{I}[T_{i}=T_{j},K_{i}=K_{j}]}{\sum_{i=1}^{n}\mathbb{I}[K_{i}=K_{j}]}$ (D.5)
@@
$\text{Risk}_{j}=\max\left\{0,\frac{TCAP^{\prime}_{j}-WEAP_{j}}{1-WEAP_{j}}\right\}$ (D.6)
A value near 0 indicates synthetic data provides no additional risk beyond baseline, while a value near $1$ signals a maximum increase in disclosure risk. Negative values (adversary performs worse than random) are truncated to zero by convention (Ran et al. 2024; Elliot et al. 2023).
Intuition: This metric isolates the extra “leakage” in matching risk—beyond what already exists in the univariate distribution of the target — arising from access to the synthetic data. It thus accounts for singling out, linkage and inference risks (due to the empirical distribution of sensitive values per equivalence class) identified by  Article 29 Data Protection Working Party (2025) as being critical determinants of risks to anonymity. The disclosure risk for the dataset is the mean of $R_{j}$ over all synthetic records meeting the $\tau$ threshold. The pseudocode to calculate TCAP can be seen in Algorithm 7
@@
11:  Calculate $\mathrm{TCAP}^{\prime}_{j}$ using equation D.4
12:  Calculate $\mathrm{WEAP}_{j}$ using equation D.5
13:  Calculate $\text{Risk}_{j}$ using equation D.6
@@
D.3 Additional Diagnostic Metrics
To complement Utility and Disclosure Risk, we report standard diagnostic metrics that quantify low-order distributional fidelity, detectability, and sample-wise quality. These diagnostics are provided for completeness and to facilitate comparison with prior synthetic tabular data benchmarks. Unless stated otherwise, lower values indicate closer agreement between the real and synthetic distributions.
D.3.1 Discrepancy between Synthetic and Real Data Distributions
We measure the discrepancy between the synthetic and real distributions using the Wasserstein distance (WD) and Maximum Mean Discrepancy (MMD). In particular, mixed-type tabular records are first mapped into a common vector space via the same preprocessing applied by the data loader (continuous normalisation and categorical encoding), yielding vectors $x\in\mathbb{R}^{d}$. On large datasets, computing the full distance is expensive due to pairwise computations between samples. Study conducted by Jolicoeur-Martineau et al. (2024) also evaluated using these measure on data with less than 5,000 rows. We therefore use a batched evaluation protocol: we draw $K$ non-overlapping (or randomly sampled) batches of size n=5,000n=5{,}000 from the real and synthetic datasets and report the average distance as seen in equation D.8.
$\widehat{W}(P_{r},P_{g})$ $=\frac{1}{K}\sum_{k=1}^{K}W\!\left(\mathcal{B}_{r}^{(k)},\mathcal{B}_{g}^{(k)}\right),$ (D.7)
$\mathrm{MMD}^{2}(P_{r},P_{g})$ $=\mathbb{E}_{x,x^{\prime}\sim P_{r}}\!\left[k(x,x^{\prime})\right]+\mathbb{E}_{y,y^{\prime}\sim P_{g}}\!\left[k(y,y^{\prime})\right]-2\,\mathbb{E}_{x\sim P_{r},y\sim P_{g}}\!\left[k(x,y)\right].$ (D.8)
where $\mathcal{B}_{r}^{(k)}$ and $\mathcal{B}_{g}^{(k)}$ are the $k$-th batches of the preprocessed real and synthetic samples, respectively. This provides a practical Monte Carlo estimate of the distributional discrepancy under limited compute.
D.3.2 Shape score (column-wise density discrepancy)
We summarisecolumn-wise density discrepancies using a shape score computed as the average of (i) the Kolmogorov–Smirnov statistic (KST) over continuous columns and (ii) TVD over categorical columns. For $j$-th continuous variable:
$\mathrm{KST}_{j}\;=\;\sup_{x}\left|F_{j}(x)-F^{\prime}_{j}(x)\right|.$ (D.9)
For $k$-th categorical variable, we use $\mathrm{TVD}(p_{k},p^{\prime}_{k})$ from equation D.10.
$\mathrm{TVD}(p_{k},p^{\prime}_{k})\;=\;\frac{1}{2}\sum_{k=1}^{D_{\text{cat}}}\left|p_{k}(\omega)-p^{\prime}_{k}(\omega)\right|.$ (D.10)
The aggregate shape score is
$\mathrm{Shape}\;=\;\frac{1}{D_{\text{num}}+D_{\text{cat}}}\left(\sum_{j=1}^{D_{\text{num}}}\mathrm{KST}_{j}+\sum_{k=1}^{D_{\text{cat}}}\mathrm{TVD}(p_{k},p^{\prime}_{k})\right).$ (D.11)
D.3.3 Trend score (pairwise dependence discrepancy)
We quantify how well pairwise dependence is preserved using a trend score that combines (i) Pearson correlation discrepancies for continuous–continuous pairs and (ii) contingency table discrepancies for categorical–categorical pairs.
Continuous–continuous (Pearson).
For continuous columns $a,b\in\mathcal{D}_{\text{num}}$, let $\rho^{\mathcal{D}}_{a,b}$ and $\rho^{\mathcal{D}^{\prime}}_{a,b}$ be Pearson correlations computed on $\mathcal{D}$ and $\mathcal{D}^{\prime}$. Since $\rho\in[-1,1]$, we scale absolute differences by $1/2$ so the score lies in $[0,1]$:
$\mathrm{Trend}_{\mathrm{cont}}\;=\;\frac{1}{2|\mathcal{P}_{\mathcal{D}_{\text{num}}}|}\sum_{(a,b)\in\mathcal{P}_{\mathcal{D}_{\text{num}}}}\left|\rho^{\mathcal{D}}_{a,b}-\rho^{\mathcal{D}^{\prime}}_{a,b}\right|,$ (D.12)
where $\mathcal{P}_{\mathcal{D}_{\text{num}}}$ is the set of unordered continuous pairs.
Categorical–categorical (contingency similarity).
For categorical columns $u,v\in\mathcal{D}_{\text{cat}}$, let $R_{u,v}(\alpha,\beta)$ and $S_{u,v}(\alpha,\beta)$ denote the empirical joint frequencies in the contingency tables on $\mathcal{D}$ and $\mathcal{D}^{\prime}$. We compute a TVD-style discrepancy:
$\mathrm{Trend}_{\mathrm{cat}}\;=\;\frac{1}{2|\mathcal{P}_{\mathcal{D}_{\text{cat}}}|}\sum_{(u,v)\in\mathcal{P}_{\mathcal{D}_{\text{cat}}}}\sum_{\alpha\in\Omega_{u}}\sum_{\beta\in\Omega_{v}}\left|R_{u,v}(\alpha,\beta)-S_{u,v}(\alpha,\beta)\right|,$ (D.13)
where $\mathcal{P}_{\mathcal{D}_{\text{cat}}}$ is the set of unordered categorical pairs.
When both continuous and categorical pairs are present, we report overall trend score as their average.
D.3.4 Detection score via classifier two-sample test (C2ST)
We measure detectability by training a binary classifier to distinguish real from synthetic records. We construct a labeled dataset by assigning label 1 to samples from $\mathcal{D}$ and label 0 to samples from $\mathcal{D}^{\prime}$, then evaluate the classifier on a held-out split. The C2ST score is the test accuracy $\mathrm{Acc}\in[0,1]$, where values close to $0.5$ indicate low detectability. Optionally, we report a symmetric detection score $\mathrm{Det}=2|\mathrm{Acc}-0.5|\in[0,1]$.
D.3.5 Sample-wise fidelity and coverage: $\alpha$-precision and $\beta$-recall
We report $\alpha$-precision and $\beta$-recall, which summarise sample-wise fidelity and coverage of the real distribution. Informally, $\alpha$-precision is the fraction of synthetic samples that fall within the “typical” (high-density) region of the real distribution, while $\beta$-recall is the fraction of real samples covered by the typical region of the synthetic distribution. We compute these scores using the procedure in, based on minimum-volume (high-density) sets estimated from samples.
D.4 Reproducibility
@@
This appendix provides the procedural summary sufficient for manual replication or verification by independent researchers. For computational resource, we run all experiments using Python 3.11 with library torch 2.3.0 in an L40S GPU. However, since the model used in this study is an MLP, it is also possible to reproduce this experiment under regular GPU like GTX/RTX family. Table 68 showed detailed variables implemented in the utility and risk evaluation.
@@
Appendix E Additional and Detailed Results
E.1 Results on Standard Tabular Generative Model Evaluations
Tables 10-13 report diagnostic metrics that complement the main Utility and Disclosure Risk scores. For WD, MMD, Shape, and Trend, lower is better as these metrics quantify discrepancies between real and synthetic statistics. For the C2ST detection score in Table 13, we follow the convention used in our evaluation code and report the detection score as defined there, where larger values correspond to stronger indistinguishability between real and synthetic samples.
Discrepancy between data distributions (WD and MMD).
Tables 9 and  10 disseminated the results on MMD and WD. From Table 9, it seems that TabSyn dominates in two datasets (Fiji, Rwanda), while in Adult and Churn results showed similar MMD values in four algorithms. Moreover, the difference of the distance of most algorithms, except TabDDPM are around 5 decimals. Therefore, MMD comparison results are inconclusive for this study, making WD becomes supporting results.
Table 9: Average MMD across 7 datasets. We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table highlights the TabSyn are best in two datasets, while in Adult and Churn most algorithms are competitive, making it hard to conclude.
Data TabDDPM TabSyn TabSynFlow TabSynFlow TabbyFlow TabbyFlow
OT VP OT VP
UK $\bm{4.0127\times 10^{-4}}$ $\underline{4.0197\times 10^{-4}}$ $4.0229\times 10^{-4}$ $4.0804\times 10^{-4}$ $4.0837\times 10^{-4}$ $4.1794\times 10^{-4}$
Canada $2.1109\times 10^{-2}$ $5.7652\times 10^{-4}$ $\underline{5.6669\times 10^{-4}}$ $\bm{5.5739\times 10^{-4}}$ $5.9633\times 10^{-4}$ $5.9156\times 10^{-4}$
Fiji $3.7338\times 10^{-1}$ $\bm{4.0162\times 10^{-4}}$ $4.0283\times 10^{-4}$ $4.0578\times 10^{-4}$ $\underline{4.0257\times 10^{-4}}$ $4.0270\times 10^{-4}$
Rwanda $2.5037\times 10^{-1}$ $\bm{4.7202\times 10^{-4}}$ $4.9067\times 10^{-4}$ $5.5161\times 10^{-4}$ $4.9030\times 10^{-4}$ $5.0168\times 10^{-4}$
Indonesia $4.0213\times 10^{-4}$ $\underline{4.0174\times 10^{-4}}$ $\bm{4.0112\times 10^{-4}}$ $4.3640\times 10^{-4}$ $4.7961\times 10^{-4}$ $4.8380\times 10^{-4}$
Adult $4.0026\times 10^{-4}$ $\bm{4.0005\times 10^{-4}}$ $\bm{4.0005\times 10^{-4}}$ $\bm{4.0005\times 10^{-4}}$ $\underline{4.0006\times 10^{-4}}$ $\bm{4.0005\times 10^{-4}}$
Churn $4.0004\times 10^{-4}$ $\bm{4.0000\times 10^{-4}}$ $\bm{4.0000\times 10^{-4}}$ $\bm{4.0000\times 10^{-4}}$ $\underline{4.0001\times 10^{-4}}$ $\bm{4.0000\times 10^{-4}}$
In contrast, Table 10 shows that TabbyFlow-OT achieves the lowest average per-continuous feature WD on five out of seven datasets (Canada, Fiji, Rwanda, Indonesia, and Adult), indicating the most faithful preservation of distributions in those settings. In contrast, TabDDPM is best on UK, and Churn. A consistent pattern is that TabbyFlow variants have lower WD values than TabSynFlow-OT across all datasets, suggesting that overall TabbyFlow could preserve the data distributions better in these benchmarks.
Table 10: Average Wasserstein Distance across 7 datasets. We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table highlights the domination of TabbyFlow as the best performers, followed by TabDDPM in three datasets, meaning that TabbyFlow is best in preserving continuous variables.
Data TabDDPM TabSyn TabSynFlow TabSynFlow TabbyFlow TabbyFlow
OT VP OT VP
UK 1.7090±0.0024 1.7680±0.0026 1.7690±0.0023 1.8622±0.0028 1.7177±0.0021 1.7675±0.0021
Canada 8.9481±0.0179 2.5620±0.0061 2.5927±0.0047 2.7237±0.0052 2.3748±0.0053 2.4522±0.0072
Fiji 9.2802±0.0100 2.5594±0.0035 2.5869±0.0038 2.7487±0.0039 2.4620±0.0032 2.5274±0.0024
Rwanda 5.9976±0.0244 0.7335±0.0029 0.7483±0.0031 0.7854±0.0022 0.6797±0.0030 0.7154±0.0032
Indonesia 0.1732±0.0007 0.1757±0.0007 0.1776±0.0006 0.1859±0.0005 0.1731±0.0006 0.1778±0.0006
Adult 0.8392±0.0024 0.7976±0.0024 0.8141±0.0023 0.8558±0.0025 0.7665±0.0023 0.7937±0.0026
Churn 0.1528±0.0027 0.1592±0.0021 0.1600±0.0020 0.1930±0.0031 0.1653±0.0023 0.1652±0.0022
Mixed-type univariate fidelity (Shape).
The Shape metric aggregates column-wise discrepancies across both continuous and categorical attributes. Table 11 indicates a split regime: TabbyFlow is strongest on UK, Canada, Fiji, and Rwanda, which aligns with its strong categorical TVD performance and its ability to preserve mixed-type univariate structure. TabSynFlow-OT is strongest on Indonesia, Adult, and Churn, which aligns with its consistently low WD and indicates that these datasets benefit more from improvements in continuous modeling or from the specific trajectory specification used by TabSynFlow-OT.
Table 11: Performance comparison on the error rates (%) of Shape. We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table highlights the domination of TabbyFlow and TabSynFlow as the best performers, meaning that the algorithms are best in preserving the distribution of mixed-type data.
Data TabDDPM TabSyn TabSynFlow TabSynFlow TabbyFlow TabbyFlow
OT VP OT VP